PCl3 + Cl2 ⇌ PCl5, calculate all three equilibrium concentrations when Kc = 16.0 and [PCl5] i = 1.00 M.
1
Expert's answer
2020-12-21T03:53:19-0500

For the reaction:

PCl3 + Cl2 ⇌ PCl5


the equilibrium concentration constant "K_c" has the following expression:

"K_c = \frac{[PCl_5]}{[PCl_3][Cl_2]}" ,

where "[PCl_5], [PCl_3]" and "[Cl_2]" are the equilibrium concentrations of PCl5, PCl3 and Cl2, respectively.


The equilibrium concentration of PCl5 is its initial concentration, 1.00 M minus the quantity of PCl5 dissociated "x" :


"[PCl_5] = 1.00 - x" .


Therefore, the equilibrium concentrations of PCl3 and Cl2 are "x" :

"[PCl_3] = [Cl_2] = x" .

Using these equations, let's find "x" :

"16.0 = \frac{1-x}{x^2}"

"16x^2 + x - 1 =0"

"x = 0.221" .


Finally, the equilibrium concentrations are:

"[PCl_5] = 1 - 0.221 = 0.779" M

"[PCl_3] = [Cl_2] = 0.221" M.


Answer: [PCl5] = 0.779 M, [PCl3] = [Cl2] = 0.221 M.